Use factoring to solve the equation
6x^2 + 23x - 18 = 0
I know the choices are
x=2/3, -2/9
x=3/2, -2/23
x=2/3, -23/2
x=3/2, -9/2
x=2/3, -9/2
Can someone please show me how to work this problem? Thanks

Answer by stanbon(57412)

Use factoring to solve the equation
6x^2 + 23x - 18 = 0
6x^2 + 27x-4x -18 = 0
3x(2x+9)-2(2x+9) = 0
(2x+9)(3x-2) = 0
---------------
x =-9/2 or x = 2/3
=============================
Cheers,
Stan H.
=============================
I know the choices are
x=2/3, -2/9
x=3/2, -2/23
x=2/3, -23/2
x=3/2, -9/2
x=2/3, -9/2